OpenStudy (anonymous):

ind all of the solutions for the following equation. tan(x) = 4 tan (x) - (sqrt)(3)

OpenStudy (anonymous):

\[\tan(x)=4\tan(x)-\sqrt{3}\]

OpenStudy (tkhunny):

Have you considered solving \(3\tan(x) = \sqrt{3}\)?

OpenStudy (anonymous):

I have not. is that the thing i need?

OpenStudy (tkhunny):

If you subtract \(\tan(x)\) from both sides and add \(\sqrt{3}\) to both sides, that is where you end up. What's next?

OpenStudy (anonymous):

Uhhh. divide each side by 3.

OpenStudy (tkhunny):

Right. Where does that lead us?

OpenStudy (anonymous):

\[\tan(x)=\frac{ \sqrt{3} }{ 3 }\]

OpenStudy (tkhunny):

...or \(\dfrac{1}{\sqrt{3}}\)

OpenStudy (anonymous):

here's the question with answers: so 1/root3. is that the solution then?? bc that's equal to a certain degreee, right?

OpenStudy (tkhunny):

We have \(\tan(x) = \dfrac{1}{\sqrt{3}}\) We still need \(x = ??\).

OpenStudy (anonymous):

inverse tangent of 1/root3

OpenStudy (anonymous):

x = 30

OpenStudy (anonymous):

now how do i know if it's 30 + 180 or 30 + 360?

OpenStudy (tkhunny):

\(... + k\cdot 180º\)

OpenStudy (anonymous):

lol. ok then. thank you very very much :)